3. A total of 100 ml of2.1 M weak acid is mixed with 300 ml of 0.1 MKoH. The pH of the final solution is 4.13. What isthe K and pKaof the weak acid? Show detailed steps. (0 pts)
Answer #1

millimoles of weak acid = 100 x 2.1 = 210

millimoles of base added = 300 x 0.1 = 30

210 - 30 = 180 millimoles acid left

[acid] = 180 / 400 = 0.45 M

[salt] = 30 / 400 = 0.075 M

pH = pKa + log [salt] /[acid]

4.13 = pKa + log 0.075 / 0.45

4.13 = pKa -0.78

pKa = 4.91

Ka = 10-pKa   = 10-4.91

Ka = 1.23 x 10-5